# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v3

@@ IV Towards a Unified Framework
Previous Section II and Section III illustrate two distinct approaches for controlling bipedal robots. Whether it is end-to-end or hierarchical, their ultimate goal is to adeptly navigate all types of terrains Both these two finally towards to the unified frameworks, can be named as Humanoid Foundation Model (HFM) and Multi-layered Adaptive Models (MLAM), which are expanded based on end-to-end, and hierarchical approaches.
In this section, we explore the conceptual design of HFM and MLAM, analyse the existing gaps and challenges in achieving these unified frameworks, and discuss recent advancements in quadruped robotics that offer valuable insights for the future development of bipedal locomotion.
IV-A Humanoid Foundation Models
Inspired by the Robot Foundation Models (RFMs) [60, 61], built on an end-to-end framework, enhance versatility by learning shared representations across different skills, HFM focus specifically on humanoid robots, which inherit advantages of RFMs but possess a higher number of DoFs and a human-like morphology, requiring more advanced motion control and adaptability. As foundation model, this structure bypasses the need for task decomposition, allowing the system to learn through large datasets and generalise across diverse locomotion tasks.
HFM with large parameters are expected to be pre-trained on large-scale robotic datasets across various platforms and later fine-tuned using DRL to adapt to specific locomotive tasks. This enables humanoid robots to continuously refine their locomotion strategies and adapt to dynamic environments through interaction. For instance, AdA [62] demonstrates a framework where an agent is first pre-trained on extensive datasets and subsequently fine-tuned with RL, allowing it to optimise policy performance and achieve near-optimal control.
Other foundation model work also provide thoughts about how to build up HFM. For example Current research on RFMs mainly about Vision-Language-Action (VLA) model on robotic arm platform, derived from VLM, VLA models directly map visual information to robot actions, enabling end-to-end decision-making without explicit intermediate representations. Recent work, such as RT-2 [63], expands upon the transformer model-based RT-1 [64], a vision-language model designed for robotic control. In RT-2, the entire model is trained using a combination of a VLM and robotic demonstrations. The actions need to be quantized as inputs and subsequently dequantized.
Challenges: The challenges of RMFs can be derived from two aspects: (1) training datasets and (2) model inference time.
  1. 1.
    Training datasets: In terms of datasets, unlike Large Language Models (LLMs) or VLMs, HMF takes multimodal inputs with heterogeneous data formats that are not inherently aligned. While LLMs process textual prompts as both input and output, and VLMs typically map images and text bidirectionally, HFM for bipedal robot must handle diverse sensory data, including vision, proprioception, motion trajectories, and object interactions, making input representation and model alignment more complex. Which lead to input representation and feature alignment challenge, which means different data modalities have different feature distributions and dimensionalities, making direct fusion difficult.
    Beside of directly making the training process difficult and high requirements of model construction, another challenge in datasets is datasets collection. Current datasets provide human motions for bipedal humanoid robot to imitate. For example, AMASS [65] consists solely of human motion capture data, while Motion-X [66] extends this by incorporating textual prompt labels corresponding to the motions. Although various references exist for robots to learn different skills, most datasets are not directly tailored for robots and require additional mapping to specific robotic platforms.
  2. 2.
    Model inference time: Bipedal robots require high-frequency action strategies to maintain stability, but real-time decision-making in HFM faces several limitations. First, RMFs using LLMs as backbones are computationally heavy, leading to slow inference due to expensive forward propagation. Second, autoregressive token generation introduces sequential dependencies, preventing parallel computation and further increasing latency. Third, onboard processors lack sufficient memory and computational power to run large models efficiently, making real-time execution challenging without model compression or hardware acceleration.
IV-B Multi-layered Adaptive Model
Unlike HFM, the MLAM follows a hierarchical framework, decomposing tasks into multiple sub-tasks. Each layer is responsible for a specific level of decision-making, with higher layers handling HL planning and lower layers executing fine-grained control. This structured approach enables the system to adapt dynamically and ultimately achieve the overall goal. Each module in the hierarchical model can be fine-tuned, replaced, or upgraded without requiring re-training of the entire model.
The discrete model is structured into three hierarchical levels: strategy, planning, and control, each progressively refining task execution. Section III has demonstrated how DRL-based policies contribute to both the HL planner and LL controller. At the highest level, the strategy layer serves as the reasoning component, enabling the robot to comprehend tasks, infer objectives, and decompose long-term goal-based tasks into structured sub-tasks, which are then passed down for execution. For example, [67] employs an LLM as a reasoning layer to interpret user commands, break them into sequential task steps, and translate them into executable code for a quadruped robot.
Challenge: In the previous Section LABEL:Sec:learned_heirarchical_framework, the challenges of the hierarchical framework were briefly discussed. However, when considering the unified framework—MLAM, additional challenges emerge that require further exploration. These challenges are:
  1. 1.
    Limited Visual Perception and Multi-Modal Integration: MLAM relies on real-time perception and decision-making, yet bipedal locomotion demands low-latency, high-precision sensory processing (e.g., vision, IMU, force sensors). Integrating these multi-modal inputs across different layers of the hierarchy remains a significant challenge. Each modality operates at different timescales and levels of abstraction, making seamless coordination and alignment difficult. Moreover, existing methods often fail to effectively capture cross-modal dependencies, limiting their applicability in dynamic scenarios.
  2. 2.
    Computational Efficiency and Real-Time Execution: MLAM must balance the quality of HL decision-making with the strict low-latency requirements of real-time control. This is especially problematic for HL reasoning models, such as LLMs, which are inherently designed for step-by-step reasoning and rely on autoregressive decoding. Generating decisions token-by-token significantly increases inference time, making these models unsuitable for tasks that demand rapid response times. These delays can lead to unstable behaviours, particularly in tasks requiring fast adaptation and precise coordination in dynamic environments.
These challenges highlight the complexity of implementing MLAM and underscore the need for innovative solutions to address these limitations in real-world applications.
@@ IV-C Fundamental challenges
The development of unified frameworks such as HFM and MLAM aims to address the challenges of bipedal locomotion in dynamic and unstructured environments. However, achieving such frameworks requires overcoming several fundamental obstacles. Among these, balancing generalisation and precision, as well as designing streamlined frameworks capable of handling complex tasks, represent two critical challenges. These challenges are explored in the following subsections.
IV-C1 Generalisation versus precision
The challenge of balancing generalisation and precision is a fundamental obstacle in achieving the unified frameworks of HFM and MLAM. For HFM, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments. For MLAM, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
@@ IV-C1 Generalisation versus precision
For HFM, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks. In contrast, MLAM struggles to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
@@
These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between HFM’s large-scale adaptability and MLAM’s hierarchical precision.
@@ IV-C2 Streamlined frameworks capable of handling complex tasks
One of the critical challenges in unifying HFM and MLAM is designing frameworks that are both simple and capable of handling complex tasks. While streamlined frameworks reduce computational overhead and potential failure points, they may lack the flexibility to adapt to diverse and dynamic environments. On the other hand, complex frameworks with multiple components, such as foot planners or terrain-specific modules [35, 36], often introduce additional computational burden and coordination challenges.
For HFM, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity. For MLAM, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency. Achieving a balance between simplicity and capability is essential for creating robust, unified frameworks.
@@
Figure 4: Towards a Unified Framework: the transition from the present frameworks (end-to-end and hierarchical frameworks) to a unified framework capable of handling diverse locomotion tasks. The present state is characterised by limitations such as fine-tuning requirements, limited gait patterns, restricted terrain walking, and blind locomotion. To bridge this gap, future developments must focus on generalisation, precision, streamlined frameworks, vision-based learning, and zero-shot sim-to-real capabilities. These advancements will lead to the creation of two potential unified models: (1) a multi-layered adaptive model grounded in the hierarchical framework and (2) a large robot model built upon the end-to-end framework. The unified framework will ultimately support diverse locomotion tasks in complex environments.
@@ V-A3 Integrating foundation models
Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots as discussed in Section IV. Their integration into robotics extends beyond decision-making and offers new ways to enhance control, adaptability, and learning efficiency.
HL task planning: Foundation models, particularly LLMs and VLMs, enable robots to process multimodal inputs, enhancing task planning and execution. By integrating visual and textual data, these models enhance context-awareness, allowing robots to make more informed, adaptive decisions.
For instance, Vision-Language Model Predictive Control (VLM-PC) [82] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. Similarly, in hierarchical frameworks, VLMs can act as an intermediary layer, processing sensory data before passing structured commands to lower-level controllers.
By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution. Future research should focus on integrating LLMs with DRL, where LLMs guide HL strategy while DRL refines real-time execution, creating an adaptive, flexible control loop. This integration would enable robots to generalize across tasks, adapt dynamically, and make autonomous complex decisions.
Language-Assisted DRL: The design of effective reward functions is a fundamental challenge in DRL for bipedal robots. While periodic reward functions have been successfully applied to facilitate cyclic movements like walking [28], defining reward structures for non-periodic tasks, such as jumping, manipulation, and complex locomotion transitions, remains an open problem. These tasks require distinct success criteria beyond simple periodic movement cycles, making manual reward tuning inefficient and time-consuming. Additionally, optimising reward functions for multi-task learning and environmental adaptability is a major bottleneck, highlighting the need for adaptive, self-adjusting reward mechanisms.
@@ VI Conclusion
Each framework brings unique strengths to the unified framework. End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility needed for robots to acquire diverse locomotion skills. In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies. Together, these approaches form the basis for two potential unified frameworks: the Humanoid Robot Model from the end-to-end approach and the Multi-Layered Adaptive Model from the hierarchical approach, both designed to meet the complex demands of bipedal robot locomotion.
@@ Appendix B Training Simulation Environment
The development of DRL algorithms and sim-to-real techniques has shown the requirement for high-quality simulators. Creating a reliable simulation environment and conducting RL training is difficult. Through the literature, several simulators are available, including Isaac Gym [72], robocup3D [73], Openai Gym [74], MuJoCo [75].
OpenAI developed Gym and Gymnasium providing a simplified environment for fast testing RL algorithms with the simplified bipedal model. Also, RoboCup serves as a platform for RL research and development. On the other hand, MuJoCo and Gazebo developed by DeepMind respectively are comprehensive physics engines that support robotics research and development, along with other applications needing physics simulation. Isaac Gym, a cutting-edge physics simulation and RL framework developed by NVIDIA, is a high-performance platform for training agents in intricate and lifelike environments.

### v4

@@ IV Towards a Unified Framework
Previous Sections II and III outlined two distinct paradigms for controlling bipedal robot locomotion: end-to-end and hierarchical frameworks. While both paradigms have demonstrated significant success in addressing specific locomotion tasks, their ability to generalise across diverse robot morphologies, environmental conditions, and locomotion objectives remains limited. Consequently, this limitation motivates us to revisit the central question posed in Section I:
Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks? If not, how should such a framework be designed, and what features or components should it include?
To address this, we first explicitly define a unified framework for bipedal locomotion as a control architecture capable of supporting consistent learning, transfer, and execution across arbitrary robot morphologies, terrains, and task definitions, ultimately aiming to develop genuinely generalist locomotion policies that achieve robust zero-shot or few-shot generalisation to previously unseen scenarios, embodiments, and objectives.
Evaluating existing approaches against this definition reveals that neither the end-to-end nor hierarchical control paradigms currently meet these unified criteria. Typically, these methods require significant redesign or retraining when introduced to new morphologies, tasks, or environmental conditions, highlighting the practical gap towards achieving genuine unification.
To explore pathways towards unified frameworks, we identify two currently foreseeable and potentially viable directions that build upon the structural foundations of the previously discussed paradigms—end-to-end and hierarchical. These two directions offer distinct but complementary strategies for advancing towards scalable, generalisable locomotion architectures.
  • Bipedal Foundation Models (BFMs) – large-scale, pre-trained models that map perception directly to action through representation learning. Trained on diverse data across tasks and embodiments, BFMs aim to enable generalist locomotion control by supporting rapid adaptation via fine-tuning.
  • Multi-Layer Adaptive Frameworks (MLAFs) – modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs. MLAFs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
In the following subsections, we present detailed conceptual analyses of BFMs and MLAFs, evaluate their potential to realise a unified framework, and subsequently discuss fundamental challenges inherent to existing frameworks.
IV-A Bipedal Foundation Models
Inspired by Robot Foundation Models [60, 61], BFMs are large-scale, general-purpose models tailored for bipedal locomotion. Trained on diverse datasets spanning various morphologies, terrains, and objectives, BFMs enable reasoning about novel scenarios and adaptation with minimal task-specific supervision. A second-stage DRL fine-tuning process further adapts these models to specific platforms, supporting versatile locomotion behaviours such as walking and turning. Building upon the end-to-end reinforcement learning methods introduced in Section II, which typically focus on task-specific policy training, BFMs retain RL as a core adaptation mechanism, but shift toward a unified and generalisable paradigm by introducing shared motion priors across tasks and morphologies. This reuse of RL in a fine-tuning role allows BFMs to ground HL priors into executable LL policies, while improving sample efficiency and transferability.
The BFMs architecture includes a multi-modal embedding module, a latent backbone, and an action decoder. The embedding encodes morphology, task specs, and optional environment input into a shared space. The backbone converts these into a temporally structured latent representation conditioned on the task. The decoder outputs LL control commands adapted to the robot. Similar modular structures appear in RT-1 [62] and RT-2 [63], which map multimodal inputs to discrete actions using a shared transformer backbone.
The model first leverages IL to pre-train generalisable motion priors from diverse demonstrations, resulting in latent representations that facilitate scenario reasoning and transfer across different robot morphologies. DRL then fine-tunes these priors by aligning them with the physical constraints, action spaces, and dynamics of specific platforms. Rather than learning from scratch, DRL enables efficient adaptation to new tasks. Lightweight strategies such as ControlNet-style modular layers such as AdaptNet [64]—allow task-specific adjustments with minimal parameter updates.
Beyond modular adaptation, recent works have demonstrated that large-scale DRL fine-tuning significantly enhances generalisation and capability transfer. FLaRe [65] fine-tunes behaviour-cloned policies with sparse rewards to achieve high success rates on long-horizon mobile manipulation tasks, including generalisation to novel embodiments. Complementarily, MOTO [66] applies a model-based DRL framework that incorporates on-policy value expansion and epistemic uncertainty estimation, enabling effective offline-to-online adaptation from image observations.
Other frameworks further highlight this paradigm. AdA [67] demonstrates that scaling DRL with memory-based policies allows for in-context adaptation to novel embodied environments. RoboFUME [68] combines VLM-based reward classification with DRL fine-tuning to autonomously adapt pre-trained manipulation policies without environment resets. Collectively, these approaches underscore DRL not only as a fine-tuning tool, but as a central mechanism for grounding foundation model priors into executable, platform-specific control policies.
While DRL enables flexible and powerful adaptation of pre-trained motion priors, applying it in bipedal foundation models introduces several challenges. Fine-tuning often demands carefully shaped rewards, frequent environment resets, and large amounts of interaction data—factors that are especially costly and risky on physical bipedal robots [68]. Additionally, DRL policies may overfit to narrow dynamics or forget generalisable priors acquired during pre-training [65, 67]. These issues constrain the robustness, efficiency, and safety of DRL-based adaptation in real-world settings. Meanwhile, existing datasets such as AMASS [69] and Motion-X [70] remain human-centric and require substantial retargeting effort [25]. The inference latency of large models also limits real-time deployment on hardware-constrained platforms [63].
IV-B Multi-layered Adaptive Models
Unlike BFMs, which rely on large-scale pre-training and shared latent representations, MLAMs adopt a modular hierarchical approach, extending conventional hierarchical frameworks in Section III with explicitly adaptive layers. MLAMs decompose the locomotion pipeline into three distinct adaptive tiers: (1) HL reasoning for task decomposition and goal interpretation, (2) mid-level planning for behaviour generation, and (3) LL control for precise command execution. This modularity allows each layer to be independently optimised, upgraded, or replaced without retraining the entire system. MLAMs distinguish themselves from conventional hierarchical approaches through their adaptive modularity, dynamic policy integration, and clear layer-wise semantic responsibilities.
A key distinguishing feature of MLAMs lies in its capacity to dynamically compose adaptive modules for each control tier. Instead of generalising through unified embeddings as in BFMs, MLAMs achieve scalability by dynamically assembling specialised policies. Each layer processes context-specific inputs and outputs interpretable commands. The HL reasoning layer typically uses symbolic reasoning or Large Language Models (LLMs) [71, 72] to parse commands into sub-tasks. For instance, Vision-Language Model Predictive Control [73] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution. The mid-level planner selects or synthesises motions via learned motion libraries [55, 41] or DRL-based planners [74].
The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands. These include locomotion primitives like walking and climbing [75], adaptive tracking controllers for whole-body tracking [56], and imitation-based skills such as kicking and dancing [22]. Controllers at this level utilise RL, IL, or model-based methods [9, 76] and operate at high frequency to ensure robust real-time execution across varied conditions. An example of MLAMs architecture is demonstrated by recent hierarchical systems using LLMs for quadrupedal tasks [75]. HL modules interpret instructions, mid-level planners select motion primitives, and LL controllers execute precise actions, enabling robust dynamic replanning and flexible real-world operation.
Despite these strengths, MLAMs deployment faces challenges distinct from BFMs. MLAMs depend on real-time multi-modal perception such as vision, IMU, force sensors, complicating data alignment across layers due to differing timescales and abstraction levels [76]. Additionally, HL reasoning modules like LLMs often incur computational latency from autoregressive inference, posing limitations for tasks needing rapid reactions [75]. Thus, integrating multi-modal perception and ensuring real-time execution are primary challenges for effective MLAMs implementation.
Taken together, MLAMs offer a complementary path to BFMs by emphasising modularity, interpretability, and adaptability through dynamic policy composition. While they avoid the scalability bottlenecks of monolithic models, their reliance on real-time multi-modal processing and multi-layer coordination introduces unique system-level challenges.
@@ IV-C Fundamental challenges
The development of unified frameworks such as BFMs and MLAMs aim to address the challenges of bipedal locomotion in dynamic and unstructured environments. However, achieving such frameworks requires overcoming several fundamental obstacles. Among these, balancing generalisation and precision, as well as designing streamlined frameworks capable of handling complex tasks, represent two critical challenges. These challenges are explored in the following subsections.
IV-C1 Generalisation and precision
The challenge of simultaneously improving generalisation across diverse skills and precision in specific tasks, such as jumping precisely to a desired distance, remains a fundamental obstacle in realising unified frameworks like BFMs and MLAMs. For BFMs, which focuses on large-scale, versatile humanoid models, the difficulty lies in enabling the model to adapt to diverse tasks while maintaining precise control in dynamic environments. For MLAMs, which emphasizes task decomposition and hierarchical control, ensuring precision across sub-tasks while maintaining a HL general policy is equally challenging.
@@ IV-C1 Generalisation and precision
For BFMs, these challenges are exacerbated by its reliance on diverse and large-scale datasets, which often lack the precision required for specific bipedal tasks. In contrast, MLAMs struggle to achieve precision across its hierarchical layers, where HL policies must generalise across tasks while LL controllers maintain stability and accuracy in dynamic scenarios.
@@
These challenges underscore the need for innovative solutions that balance task generalisation with control precision, bridging the gap between BFM’s large-scale adaptability and MLAM’s hierarchical precision.
@@ IV-C2 Streamlined frameworks capable of handling complex tasks
One of the critical challenges in unifying BFMs and MLAMs is designing frameworks that are both simple and capable of handling complex tasks. While streamlined frameworks reduce computational overhead and potential failure points, they may lack the flexibility to adapt to diverse and dynamic environments. On the other hand, complex frameworks with multiple components, such as foot planners or terrain-specific modules [35, 36], often introduce additional computational burden and coordination challenges.
For BFMs, the challenge lies in creating a model that leverages large-scale general representations without adding unnecessary layers of complexity. For MLAMs, the difficulty is in coordinating hierarchical decision-making across layers without introducing redundancy or inefficiency. Achieving a balance between simplicity and capability is essential for creating robust, unified frameworks.
@@
Figure 4: Towards a Unified Framework: the transition from the present frameworks (end-to-end and hierarchical frameworks) to a unified framework capable of handling diverse locomotion tasks. The present state is characterised by limitations such as fine-tuning requirements, limited gait patterns, restricted terrain walking, and blind locomotion. To bridge this gap, future developments must focus on generalisation, precision, streamlined frameworks, vision-based learning, and zero-shot sim-to-real capabilities. These advancements will lead to the creation of two potential unified models: (1) a Multi-Layered Adaptive Model grounded in the hierarchical framework and (2) a Bipedal Foundation Model built upon the end-to-end framework. The unified framework will ultimately support diverse locomotion tasks in complex environments.
@@ V-A3 Integrating foundation models
Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots, as discussed in Section IV. Their integration into robotics extends beyond decision-making, offering new opportunities to enhance control, adaptability, and learning efficiency. One particularly promising direction is their use in RL, not merely for planning but for shaping the learning process itself—most notably in the design of reward functions, which remains a core challenge in bipedal locomotion tasks.
The design of effective reward functions is a fundamental challenge in DRL for bipedal robots. While periodic reward functions have been successfully applied to facilitate cyclic movements like walking [28], defining reward structures for non-periodic tasks, such as jumping, manipulation, and complex locomotion transitions, remains an open problem. These tasks require distinct success criteria beyond simple periodic movement cycles, making manual reward tuning inefficient and time-consuming. Additionally, optimising reward functions for multi-task learning and environmental adaptability is a major bottleneck, highlighting the need for adaptive, self-adjusting reward mechanisms.
@@ VI Conclusion
Each framework brings unique strengths to the unified framework. End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility needed for robots to acquire diverse locomotion skills. In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies. Together, these approaches form the basis for two potential unified frameworks: the Bipedal Robot Model from the end-to-end approach and the Multi-Layered Adaptive Model from the hierarchical approach, both designed to meet the complex demands of bipedal robot locomotion.
@@ Appendix B Training Simulation Environment
The development of DRL algorithms and sim-to-real techniques has shown the requirement for high-quality simulators. Creating a reliable simulation environment and conducting RL training is difficult. Through the literature, several simulators are available, including Isaac Gym [81], robocup3D [82], Openai Gym [83], MuJoCo [84], Orbit [117], Brax [118], Isaac Lab [119].
OpenAI developed Gym and Gymnasium to provide lightweight environments for rapid testing of RL algorithms, including simplified bipedal locomotion models. RoboCup also serves as a benchmark platform for RL research and development in multi-agent settings.
For physics-based simulation, MuJoCo developed by DeepMind, and Gazebo are widely used platforms that support a range of robotics research tasks. NVIDIA’s Isaac Gym, although now deprecated, played an important role as a high-performance GPU-based simulator for training agents in complex environments. Its successors, such as Isaac Lab and Orbit, continue to evolve as modern RL and robotics frameworks.